f(x)=tanx-1/secx


1
Expert's answer
2021-05-24T18:51:04-0400

"f(x)=\dfrac{\tan x-1}{\sec x}"

"f(x)=\dfrac{\dfrac{\sin x}{\cos x}-1}{\dfrac{1}{\cos x}}"

"f(x)=\sin x-\cos x"


Differentiating the above equation

"\dfrac{d[f(x)]}{dx}=\cos x+\sin x"

"f'(x)=\sin x+\cos x"